1. Let yi = Bo + B12; + ; where u~ N(0,02). Let û = y - Bo – Budi. Find E(u - ū)2
Answer #1

Comider the given problem here the regression exs is given by – Yi = 60 + 6, Xitu, where wir nos - nn = 4i-bo-b, xi - in = Y;